# The strong convergence phenomenon

### v1

@@ 1.1. Optimal spectral gaps
Let us begin by recalling that the proof of Lemma 1.2 is based on the simple observation that for any graph $G$, the number of closed walks with a given length and starting vertex is lower bounded by the number of such walks in its universal cover $\tilde{G}$. When $G$ is $d$-regular, its universal cover $\tilde{G}$ is the infinite $d$-regular tree. From this, it is not difficult to deduce that the maximum nontrivial eigenvalue of a $d$-regular graph is asymptotically lower bounded by the spectral radius of the infinite $d$-regular tree, which is $2\sqrt{d-1}$ [71, §5.2.2].
@@
Figure 1.1. Left figure: $4$-regular graph generated by two permutations. Right figure: $4$-regular tree generated by two free generators $a,b$ of the free group $\mathbf{F}_{2}$. The edges defined by the two generators are colored red and blue, respectively.
@@
where $U_{i}^{N}$ are i.i.d. random permutation matrices of dimension $N$ [^2] This is the permutation model of random graphs; see [50, p. 3] for its relation to other models.
@@ 1. $\bullet$
    It was shown by Greenberg [71, Theorem 6.6] that for any sequence of finite graphs $G^{N}$ with diverging number of vertices that have the same universal cover $\tilde{G}$, the maximal nontrivial eigenvalue of $G^{N}$ is asymptotically lower bounded by the spectral radius of $\tilde{G}$. On the other hand, given any (not necessarily regular) finite graph $G$, there is a natural construction of random lifts $G^{N}$ with the same universal cover $\tilde{G}$ [71, §6.1]. It was shown by Bordenave and Collins [19] that an optimal spectral gap phenomenon holds for random lifts of any graph $G$.
@@ 2. $\bullet$
Another series of developments surrrounding optimal spectral gaps arises from a different perspective on Theorem 1.4. The map $\mathrm{std}_{N}:\mathbf{S}_{N}\to\mathrm{M}_{N-1}(\mathbb{C})$ that assigns to each permutation $\sigma\in\mathbf{S}_{N}$ the restriction of the associated $N\times N$ permutation matrix to $1^{\perp}$ defines an irreducible representation of the symmetric group $\mathbf{S}_{N}$ of dimension $N-1$ called the standard representation. Thus
@@
Recent advances in the study of strong convergence have led to major progress in the understanding of such questions [21, 32, 33, 90, 30]. One of the most striking results in this direction to date is the recent work of Cassidy [30], who shows that strong convergence for the symmetric group holds uniformly for all nontrivial irreducible representations of $\mathbf{S}_{N}$ of dimension up to $\exp(N^{\frac{1}{12}-\delta})$. [^4] For comparison, all irreducible representations of $\mathbf{S}_{N}$ have dimension $\exp(O(N\log N))$. This makes it possible, for example, to study natural models of random regular graphs that achieve optimal spectral gaps using far less randomness than is required by Theorem 1.4. We will discuss these results in more detail in section 5.3.
@@ 1.2. Intrinsic freeness
[…]
@@ 1.3. New methods in random matrix theory
The development of strong convergence has gone hand in hand with the discovery of new methods in random matrix theory. For example, Haagerup and Thorbjørnsen [60] pioneered the use of self-adjoint linerization (section 2.5), which enabled them to make effective use of Schwinger-Dyson equations to capture general polynomials (their work was extended to various classical random matrix models in [121, 7, 41]); while Bordenave and Collins [19, 21, 20] developed operator-valued nonbacktracking methods in order to efficiently apply the moment method to strong convergence.
@@ 1.3.1. The interpolation method
[…]
@@ 2. $\bullet$
    A corollary of the Markov inequality states that $\|f\|_{L^{\infty}[-1,1]}\leq 2\max_{x\in I}|f(x)|$ for any discretization $I$ of $[-1,1]$ with spacing at most $\frac{1}{q^{2}}$.
@@
[…]
@@
When it is applicable, the polynomial method has typically provided the best known quantitative results and has made it possible to address previously inaccessible questions. To date, this includes works of Magee and de la Salle [90] and of Cassidy [30] on strong convergence of high dimensional representations of the unitary and symmetric groups (see also [33]); strong convergence for polynomials with coefficients in subexponential operator spaces [33]; strong convergence of the tensor GUE model of graph products of semicircular variables (ibid.); a characterization of the unusual large deviations in Friedman’s theorem [32] as illustrated in Figure 1.3; and work of Magee, Puder and the author on strong convergence of uniformly random permutation representations of surface groups [92].
@@ 2.1. $C^{*}$-probability spaces
Given a $C^{*}$-probability space $(\mathcal{A},\tau)$, we can now associate to each self-adjoint element $a\in\mathcal{A}$, $a=a^{*}$ a spectral distribution $\mu_{a}$ by defining
[…]
@@
Let $(\mathcal{A},\tau)$ be a $C^{*}$-probability space. Then for any self-adjoint $a\in\mathcal{A}$, $a=a^{*}$, we have $\mathop{\mathrm{supp}}\mu_{a}=\mathrm{sp}(a)$ and thus
@@
[…]
@@
Let $\mathbf{G}$ be a finitely generated group with generators $g_{1},\ldots,g_{r}$, and $\lambda:\mathbf{G}\to B(l^{2}(\mathbf{G}))$ be its left-regular representation, i.e., $\lambda(g)\delta_{w}=\delta_{gw}$ where $\delta_{e}\in l^{2}(\mathbf{G})$ is the coordinate vector of $w\in\mathbf{G}$. Then
@@ 2. $\bullet$
As the moments grow at most exponentially in $p$, this uniquely determines $\mu_{a}$. The density of $\mu_{a}$ was computed in a classic paper of Kesten [80, proof of Theorem 3], and is known as the Kesten distribution. Since the explicit formula for the density shows that $\mathop{\mathrm{supp}}\mu_{a}=[-2\sqrt{2r-1},2\sqrt{2r-1}]$, Lemma 2.6 yields
@@
    For every self-adjoint $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ and $f\in C(\mathbb{R})$, we have
    $\lim_{N\to\infty}\|f(P(\bm{X}^{N},\bm{X}^{N*}))\|=\|f(P(\bm{x},\bm{x}^{*}))\|\quad\text{in probability}.$
@@
    For every self-adjoint $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$, we have
    $\lim_{N\to\infty}\mathrm{d_{H}}\big{(}\mathrm{sp}(P(\bm{X}^{N},\bm{X}^{N*})),\mathrm{sp}(P(\bm{x},\bm{x}^{*}))\big{)}=0\quad\text{in probability}.$
@@
To show that $b\Rightarrow c$, we may choose for any $\varepsilon>0$ a univariate real polynomial $h$ so that $\|f-h\|_{L^{\infty}[-K,K]}\leq\varepsilon$, where $K=2\|P(\bm{x},\bm{x}^{*})\|$. Since (2.2) implies that $\|P(\bm{X}^{N},\bm{X}^{N*})\|\leq K$ with probability $1-o(1)$ as $N\to\infty$, we obtain
@@
with probability $1-o(1)$ as $N\to\infty$ by applying (2.2) to $h\circ P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$. That $b\Rightarrow c$ follows by taking $\varepsilon\downarrow 0$.
@@
with probability $1-o(1)$ as $N\to\infty$ for every $\varepsilon>0$. On the other hand, for any $y\in\mathrm{sp}(P(\bm{x},\bm{x}^{*}))$, we may choose $f\in C(\mathbb{R})$ so that $f(y)=1$ and $f(x)<1$ for $x\neq y$. Since $c$ implies that $\|f(P(\bm{X}^{N},\bm{X}^{N*}))\|\to 1$ in probability,
@@
with probability $1-o(1)$ as $N\to\infty$ for every $\varepsilon>0$. As $\mathrm{sp}(P(\bm{x},\bm{x}^{*}))$ can be covered by a finite number of such sets $y+[-\tfrac{\varepsilon}{2},\tfrac{\varepsilon}{2}]$, the implication $c\Rightarrow d$ follows. ∎
The elementary equivalent formulations of weak and strong convergence discussed above are all concerned with the (real) eigenvalues of self-adjoint polynomials. In contrast, what implications weak or strong convergence may have for the empirical distributions of the complex eigenvalues of non-self-adjoint (or non-normal) polynomials is poorly understood; see section 6.9. We nonetheless record one easy observation in this direction [100, Remark 3.6].
@@ 2.3. Strong implies weak
    For every $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$
    $\|P(\bm{X}^{N},\bm{X}^{N*})\|\leq\|P(\bm{x},\bm{x}^{*})\|+o(1)\quad\text{with probability}\quad 1-o(1).$
@@
    For every $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$
    $\|P(\bm{X}^{N},\bm{X}^{N*})\|\geq\|P(\bm{x},\bm{x}^{*})\|-o(1)\quad\text{with probability}\quad 1-o(1).$
@@
with probability $1-o(1)$ for every $p\in\mathbb{N}$, as $|P|^{2p}=(P^{*}P)^{p}\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$. The conclusion follows by letting $p\to\infty$ and applying Lemma 2.6.
@@
The implication $a\Rightarrow c$ of Lemma 2.13 also holds under the alternative hypothesis that $C^{*}(\bm{x})$ is a simple $C^{*}$-algebra; see [86, pp. 16–19].
@@ 2.4. Scalar, matrix, and operator coefficients
    $\bm{X}^{N}$ converges weakly to $\bm{x}$, i.e., for all $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$
    $\lim_{N\to\infty}\mathop{\mathrm{tr}}(P(\bm{X}^{N},\bm{X}^{N*}))=\tau(P(\bm{x},\bm{x}^{*}))\quad\text{in probability}.$
@@
    For any $C^{*}$-probability space $(\mathcal{B},\sigma)$ and $P\in\mathcal{B}\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$
    $\lim_{N\to\infty}(\sigma\otimes\mathop{\mathrm{tr}})(P(\bm{X}^{N},\bm{X}^{N*}))=(\sigma\otimes\tau)(P(\bm{x},\bm{x}^{*}))\quad\text{in probability}.$
@@
    $\bm{X}^{N}$ converges strongly to $\bm{x}$, i.e., for all $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$
    $\lim_{N\to\infty}\|P(\bm{X}^{N},\bm{X}^{N*})\|=\|P(\bm{x},\bm{x}^{*})\|\quad\text{in probability}.$
@@
    For every $D\in\mathbb{N}$ and $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$
    $\lim_{N\to\infty}\|P(\bm{X}^{N},\bm{X}^{N*})\|=\|P(\bm{x},\bm{x}^{*})\|\quad\text{in probability}.$
@@
for probability $1-o(1)$ as $N\to\infty$ for every $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$. Now note that since $\|P\|^{2p}=\|(P^{*}P)^{p}\|$ and $(P^{*}P)^{p}\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ for every $p\in\mathbb{N}$, applying the above inequality to $(P^{*}P)^{p}$ implies a fortiori that
@@
for probability $1-o(1)$ as $N\to\infty$. Taking $p\to\infty$ completes the proof. ∎
@@
A much weaker requirement is provided by the following property of $C^{*}$-algebras. We give the definition in the form that is most relevant for our purposes; its equivalence to the original more algebraic definition (in terms of short exact sequences) is a nontrivial fact due to Kirchberg, see [115, Chapter 17] or [25].
@@
with probability $1-o(1)$ as $N\to\infty$ by Lemma 2.16, while
@@
The exactness property turns out to arise frequently in practice. In particular, $C^{*}_{\rm red}(\mathbf{F}_{r})$ is exact [115, Corollary 17.10], as is $C^{*}_{\rm red}(\mathbf{G})$ for many other groups $\mathbf{G}$. For an extensive discussion, see [25, Chapter 5] or [6].
One reason that exactness is very useful in a strong convergence context is that it enables us construct complex strong convergence models by combining simpler building blocks, as will be explained briefly in section 5.4. Another useful application of exactness is that it enables an improved form of Lemma 2.13 with uniform bounds over polynomials with matrix coefficients of any dimension [90, §5.3].

### v2

@@ 1.1. Optimal spectral gaps
Let us begin by recalling that the proof of Lemma 1.2 is based on the simple observation that for any graph $G$, the number of closed walks with a given length and starting vertex is lower bounded by the number of such walks in its universal cover $\tilde{G}$. When $G$ is $d$-regular, its universal cover $\tilde{G}$ is the infinite $d$-regular tree. From this, it is not difficult to deduce that the maximum nontrivial eigenvalue of a $d$-regular graph is asymptotically lower bounded by the spectral radius of the infinite $d$-regular tree, which is $2\sqrt{d-1}$ [72, §5.2.2].
@@
Figure $1.$ Left figure: 44-regular graph generated by two permutations. Right figure: 44-regular tree generated by two free generators $a,b$ of the free group $\mathbf{F}_{2}$. The edges defined by the two generators are colored red and blue, respectively.
@@
where $U_{i}^{N}$ are i.i.d. random permutation matrices of dimension NN. [^2] This is the permutation model of random graphs; see [50, p. 3] for its relation to other models.
@@ 1. $\bullet$
    It was shown by Greenberg [72, Theorem 6.6] that for any sequence of finite graphs $G^{N}$ with diverging number of vertices that have the same universal cover $\tilde{G}$, the maximal nontrivial eigenvalue of $G^{N}$ is asymptotically lower bounded by the spectral radius of $\tilde{G}$. On the other hand, given any (not necessarily regular) finite graph $G$, there is a natural construction of random lifts $G^{N}$ with the same universal cover $\tilde{G}$ [72, §6.1]. It was shown by Bordenave and Collins [19] that an optimal spectral gap phenomenon holds for random lifts of any graph $G$.
@@ 2. $\bullet$
Another series of developments surrrounding optimal spectral gaps arises from a different perspective on Theorem 1.4. The map $\mathrm{std}_{N}:\mathbf{S}_{N}\to\mathrm{M}_{N-1}(\mathbb{C})$ that assigns to each permutation $\sigma\in\mathbf{S}_{N}$ the restriction of the associated $N\times N$ permutation matrix to $1^{\perp}$ defines an irreducible representation of the symmetric group $\mathbf{S}_{N}$ of dimension N1N-1 called the standard representation. Thus
@@
Recent advances in the study of strong convergence have led to major progress in the understanding of such questions [21, 32, 33, 91, 30]. One of the most striking results in this direction to date is the recent work of Cassidy [30], who shows that strong convergence for the symmetric group holds uniformly for all nontrivial irreducible representations of $\mathbf{S}_{N}$ of dimension up to $\exp(N^{\frac{1}{20}-\delta})$. [^4] For comparison, all irreducible representations of $\mathbf{S}_{N}$ have dimension $\exp(O(N\log N))$. This makes it possible, for example, to study natural models of random regular graphs that achieve optimal spectral gaps using far less randomness than is required by Theorem 1.4. We will discuss these results in more detail in section 5.3.
@@ 1.2. Intrinsic freeness
$\lim_{N\to\infty}\mathrm{d_{H}}\big(\mathrm{sp}(X^{N}),\mathrm{sp}(X_{\rm free})\big)=0\quad\text{a.s.}$
@@ 1.3. New methods in random matrix theory
The development of strong convergence has gone hand in hand with the discovery of new methods in random matrix theory. For example, Haagerup and Thorbjørnsen [60] pioneered the use of self-adjoint linearization (section 2.5), which enabled them to make effective use of Schwinger-Dyson equations to capture general polynomials (their work was extended to various classical random matrix models in [123, 7, 41]); while Bordenave and Collins [19, 21, 20] developed operator-valued nonbacktracking methods in order to efficiently apply the moment method to strong convergence.
@@ 1.3.1. The interpolation method
$|\mathbf{E}[\mathop{\mathrm{tr}}h(X)]-\tau(h(X_{\rm free}))|\leq\int_{0}^{1}\bigg|\frac{d}{dt}\mathbf{E}[\mathop{\mathrm{tr}}h(X(t))]\bigg|\,dt,$
@@ 2. $\bullet$
    A corollary of the Markov inequality states that $\|f\|_{L^{\infty}[-1,1]}\leq 2\max_{x\in I}|f(x)|$ for any discretization $I$ of [1,1][-1,1] with spacing at most $\frac{1}{q^{2}}$.
@@
$\scriptstyle x$$\scriptstyle I(x)$$\scriptstyle 2\sqrt{d-1}$$\scriptstyle\rho_{m_{*}+1}$$\scriptstyle\rho_{m_{*}+2}$$\scriptstyle\rho_{m_{*}+3}$$\scriptstyle\rho_{d/2-1}$$\scriptstyle d$$\scriptstyle 0$$\scriptstyle-m_{*}$$\scriptstyle-m_{*}-1$$\scriptstyle-m_{*}-2$$\scriptstyle-d/2+1\scriptstyle$-$\infty\cdots$
@@
When it is applicable, the polynomial method has typically provided the best known quantitative results and has made it possible to address previously inaccessible questions. To date, this includes works of Magee and de la Salle [91] and of Cassidy [30] on strong convergence of high dimensional representations of the unitary and symmetric groups (see also [33]); strong convergence for polynomials with coefficients in subexponential operator spaces [33]; strong convergence of the tensor GUE model of graph products of semicircular variables (ibid.); a characterization of the unusual large deviations in Friedman’s theorem [32] as illustrated in Figure 1.3; work of Magee, Puder and the author on strong convergence of uniformly random permutation representations of surface groups [94]; and work of Hide–Macera–Thomas on spectral gaps of Weil–Petersson random surfaces [69].
@@ 2.1. $C^{*}$-probability spaces
Given a $C^{*}$-probability space $(\mathcal{A},\tau)$, we can now associate to each self-adjoint element $\in\mathcal{A}, a=a$$a^{*}$ a spectral distribution $\mu_{a}$ by defining
$\int f\,d\mu_{a}=\tau(f(a))$
@@
Let $(\mathcal{A},\tau)$ be a $C^{*}$-probability space. Then for any self-adjoint $\in\mathcal{A}, a=a$$a^{*}$, we have $\mathop{\mathrm{supp}}\mu_{a}=\mathrm{sp}(a)$ and thus
@@
By the definition of support, $x\not\in\mathop{\mathrm{supp}}\mu_{a}$ if and only if there is a continuous nonnegative function $f$ so that $f(x)>0$ and $\int f\,d\mu_{a}=0$. On the other hand, by the spectral theorem, $x\not\in\mathrm{sp}(a)$ if and only if there is a continuous nonnegative function $f$ so that $f(x)>0$ and $f(a)=0$. That $\mathop{\mathrm{supp}}\mu_{a}=\mathrm{sp}(a)$ therefore follows as $\tau(f(a))=0$ if and only if $f(a)=0$, since $\tau$ is faithful and $f\geq 0$. ∎
@@
Let $\mathbf{G}$ be a finitely generated group with generators $g_{1},\ldots,g_{r}$, and $\lambda:\mathbf{G}\to B(l^{2}(\mathbf{G}))$ be its left-regular representation, i.e., $\lambda(g)\delta_{w}=\delta_{gw}$ where $\delta_{w}\in l^{2}(\mathbf{G})$ is the coordinate vector of $w\in\mathbf{G}$. Then
@@ 2. $\bullet$
As the moments grow at most exponentially in $p$, this uniquely determines $\mu_{a}$. The density of $\mu_{a}$ was computed in a classic paper of Kesten [81, proof of Theorem 3], and is known as the Kesten distribution. Since the explicit formula for the density shows that $\mathop{\mathrm{supp}}\mu_{a}=[-2\sqrt{2r-1},2\sqrt{2r-1}]$, Lemma 2.6 yields
@@
    For every self-adjoint $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ and $f\in C(\mathbb{R})$, we have lim_N→∞ f(P(X^N,X^N*)) = f(P(x,x^*))in probability.
@@
    For every self-adjoint $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$, we have lim_N→∞ d_H(sp(P(X^N,X^N*)), sp(P(x,x^*))) = 0  in probability.
@@
To show that $b\Rightarrow c$, we may choose for any $\varepsilon>0$ a univariate real polynomial $h$ so that $\|f-h\|_{L^{\infty}[-K,K]}\leq\varepsilon$, where $K=2\|P(\bm{x},\bm{x}^{*})\|$. Since (2.2) implies that $\|P(\bm{X}^{N},\bm{X}^{N*})\|\leq K$ with probability 1o(1)1-o(1) as $N\to\infty$, we obtain
@@
with probability 1o(1)1-o(1) as $N\to\infty$ by applying (2.2) to $h\circ P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$. That $b\Rightarrow c$ follows by taking $\varepsilon\downarrow 0$.
@@
with probability 1o(1)1-o(1) as $N\to\infty$ for every $\varepsilon>0$. On the other hand, for any $y\in\mathrm{sp}(P(\bm{x},\bm{x}^{*}))$, we may choose $f\in C(\mathbb{R})$ so that $f(y)=1$ and $f(x)<1$ for $x\neq y$. Since $c$ implies that $\|f(P(\bm{X}^{N},\bm{X}^{N*}))\|\to 1$ in probability,
@@
with probability 1o(1)1-o(1) as $N\to\infty$ for every $\varepsilon>0$. As $\mathrm{sp}(P(\bm{x},\bm{x}^{*}))$ can be covered by a finite number of such sets $y+[-\tfrac{\varepsilon}{2},\tfrac{\varepsilon}{2}]$, the implication $c\Rightarrow d$ follows. ∎
The elementary equivalent formulations of weak and strong convergence discussed above are all concerned with the (real) eigenvalues of self-adjoint polynomials. In contrast, what implications weak or strong convergence may have for the empirical distributions of the complex eigenvalues of non-self-adjoint (or non-normal) polynomials is poorly understood; see section 6.9. We nonetheless record one easy observation in this direction [102, Remark 3.6].
@@ 2.3. Strong implies weak
    For every $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ P(X^N,X^N*) ≤∥P(x,x^*) + o(1)  with probability 1-o(1).
@@
    For every $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ P(X^N,X^N*) ≥∥P(x,x^*) - o(1)  with probability 1-o(1).
@@
with probability 1o(1)1-o(1) for every $p\in\mathbb{N}$, as $|P|^{2p}=(P^{*}P)^{p}\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$. The conclusion follows by letting $p\to\infty$ and applying Lemma 2.6.
@@
The implication $a\Rightarrow c$ of Lemma 2.13 also holds under the alternative hypothesis that $C^{*}(\bm{x})$ is a simple $C^{*}$-algebra; see [87, pp. 16–19].
@@ 2.4. Scalar, matrix, and operator coefficients
    $\bm{X}^{N}$ converges weakly to $\bm{x}$, i.e., for all $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ lim_N→∞ tr(P(X^N,X^N*)) = τ(P(x,x^*))  in probability.
@@
    For any $C^{*}$-probability space $(\mathcal{B},\sigma)$ and $P\in\mathcal{B}\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ lim_N→∞tr)(P(X^N,X^N*)) = (στ)(P(x,x^*))  in probability.
@@
    $\bm{X}^{N}$ converges strongly to $\bm{x}$, i.e., for all $P\in\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ lim_N→∞ P(X^N,X^N*) = P(x,x^*)in probability.
@@
    For every $D\in\mathbb{N}$ and $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ lim_N→∞ P(X^N,X^N*) = P(x,x^*)in probability.
@@
for probability 1o(1)1-o(1) as $N\to\infty$ for every $P\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$. Now note that since $\|P\|^{2p}=\|(P^{*}P)^{p}\|$ and $(P^{*}P)^{p}\in\mathrm{M}_{D}(\mathbb{C})\otimes\mathbb{C}\langle x_{1},\ldots,x_{2r}\rangle$ for every $p\in\mathbb{N}$, applying the above inequality to $(P^{*}P)^{p}$ implies a fortiori that
@@
for probability 1o(1)1-o(1) as $N\to\infty$. Taking $p\to\infty$ completes the proof. ∎
@@
A much weaker requirement is provided by the following property of $C^{*}$-algebras. We give the definition in the form that is most relevant for our purposes; its equivalence to the original more algebraic definition (in terms of short exact sequences) is a nontrivial fact due to Kirchberg, see [117, Chapter 17] or [25].
@@
with probability 1o(1)1-o(1) as $N\to\infty$ by Lemma 2.16, while
@@
The exactness property turns out to arise frequently in practice. In particular, $C^{*}_{\rm red}(\mathbf{F}_{r})$ is exact [117, Corollary 17.10], as is $C^{*}_{\rm red}(\mathbf{G})$ for many other groups $\mathbf{G}$. For an extensive discussion, see [25, Chapter 5] or [6].
One reason that exactness is very useful in a strong convergence context is that it enables us construct complex strong convergence models by combining simpler building blocks, as will be explained briefly in section 5.4. Another useful application of exactness is that it enables an improved form of Lemma 2.13 with uniform bounds over polynomials with matrix coefficients of any dimension [91, §5.3].